OpenStudy (anonymous):

evaluate the expression (4x^3)^2 for x = 2

OpenStudy (anonymous):

just plug 2 in haha

OpenStudy (anonymous):

2^3 = 8

OpenStudy (anonymous):

(8*4)^2

OpenStudy (anonymous):

(32)^2 =?

OpenStudy (ibbutibbu.):

(4x^3)^2 (4(2)^3)^2 (8^3)^2 512^2 262144, i think

OpenStudy (ibbutibbu.):

no, wait nvrmind

OpenStudy (ibbutibbu.):

i screwed up

OpenStudy (anonymous):

i know I have to plug in two but it keep giving me the wrog answer so I really need you guyd help cause its timing me

OpenStudy (anonymous):

look (4x^3)^2 (4(2)^3)^2 us PEMDAS (parenthesis, exponents, multiply, divide, add, subtract) so first you use the exponent (2)^3 which ='s 8 (4(8))^2 now you have to multiply 8*4 because it's in the parenthesis (32)^2 then you do 32^2 which ='s1024

OpenStudy (anonymous):

*use

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

nooo problem